how would i solve thge system of equations using substition method?
3=5x-3y+10 &
y=2x+4 ????

Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



Your second equation is already solved in terms of y, so substitute the expression in x to which y is equal for y in the first equation, thus:




Now solve for x. Once you have a value for x, substitute that value into either of the original equations (I would suggest using the second equation because the arithmetic will be a little easier) to determine the value for y.

John


Answer by Alan3354(69443)

how would i solve thge system of equations using substition method?
3=5x-3y+10 --> 5x-3y = -7
y=2x+4
--------
Since y is already isolated in the 2nd eqn, sub for y into the 1st
5x-3y = -7
5x - 3(2x+4) = -7
5x -6x-12 = -7
-x = 5
x = -5
-----------
y = 2x+4
y = -6